OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope. A(5, 5), m = 3

sam (.sam.):

well just use the formula \[y-y_1=m(x-x_1)\] \[y-(5)=3(x-5)\] y=3x-15+5 y=3x-10